# TUMTraf EMOT: Event-Based Multi-Object Tracking Dataset and Baseline for Traffic Scenarios

### v1

@@ Abstract
In Intelligent Transportation Systems (ITS), multi-object tracking is primarily based on frame-based cameras. However, these cameras tend to perform poorly under dim lighting and high-speed motion conditions. Event cameras, characterized by low latency, high dynamic range and high temporal resolution, have considerable potential to mitigate these issues. Compared to frame-based vision, there are far fewer studies on event-based vision. To address this research gap, we introduce an initial pilot dataset tailored for event-based ITS, covering vehicle and pedestrian detection and tracking. We establish a tracking-by-detection benchmark with a specialized feature extractor based on this dataset, achieving excellent performance.
@@ I Introduction
Figure 1: This figure shows the project of TUMTraf, all the data are collected on the roadside of A9 in Garching by Munich. The second and third rows of our dataset display a sequence from left to right: the original image, a visualization of the detection and the tracking.

### v2

@@ Abstract
In Intelligent Transportation Systems (ITS), multi object tracking is mainly based on frame based cameras. However, these cameras often suffer from performance degradation under dim lighting conditions and fast motion, which are common in real world driving scenarios. Event cameras provide an alternative sensing modality with low latency, high dynamic range and high temporal resolution, making them well suited for safety critical vehicle perception tasks. Despite these advantages, existing event based vision studies in ITS are still limited in scale, and most available datasets focus on detection rather than multi object tracking. To address this gap, we introduce an initial pilot dataset designed for event based ITS applications, covering vehicle and pedestrian detection and tracking in urban traffic scenes. Based on this dataset, we provide a reference tracking by detection benchmark with an improved detector integrated into a standard tracking pipeline, aiming to serve as a baseline for future research on event based multi object tracking.
@@ I Introduction
Figure 1: This figure shows the project of TUMTraf, all the data are collected on the roadside of A9 in Garching by Munich. For more details on the dataset, visit Project Providentia++ A9 Dataset. The second and third rows of our dataset display a sequence from left to right: the original image, a visualization of the detection and the tracking.
@@ VII Conclusion
Acknowledgement
This work is supported by the MANNHEIM-CeCaS (Central Car Server – Supercomputing for Automotive, No. 16ME0820), in part by Tongji-Qomolo Autonomous Driving Commercial Vehicle Joint Lab Project, and in part by Xiaomi Young Talents Program.